OpenStudy (anonymous):

What is the equation of the tangent line to the graph of f(x)=x^4 + 2x^2 at the point where f'(x) = 1?

hero (hero):

Just use the equation for tangent line approximation

OpenStudy (anonymous):

what does that mean?

hero (hero):

y - f(c) = f'(c) (x-c) it's similar to the point slope formula

hero (hero):

But just with function notation

hero (hero):

We have to find c .... Do do this we of course have to find the derivative of f(x)

hero (hero):

Which is f'(x) = 4x^3 +4x

hero (hero):

Now set f'(x) = 1 and solve for x 1 = 4x^3 + 4x looks like you're gonna need your calculator for this one

hero (hero):

You have to find the value of x that makes this statement 1 = 4x^3 + 4x true

hero (hero):

It's complex....what kind of bizarre problem is this?

OpenStudy (anonymous):

haha. ap calculus. calculator allowed section

hero (hero):

Good luck with that

OpenStudy (anonymous):

how can I find this on the calc?

OpenStudy (amistre64):

oh that is ugly ....

hero (hero):

You won't

OpenStudy (amistre64):

you sure you aint typoed it?

OpenStudy (anonymous):

What is an equation of the line tangent to the graph of f(x) = x^4 + 2x^2 at the point where f'(x) = 1?

OpenStudy (anonymous):

haha word for word

hero (hero):

AP Calculus...gotta love it

hero (hero):

If amistre doesn't know, I'd say you're pretty much SOL

OpenStudy (anonymous):

SOL? not familiar with that abbreviation

hero (hero):

What does the back of the book say?

OpenStudy (anonymous):

It's from an actual old AP... our teacher won't give us the answers until tomorrow. SOl haha

hero (hero):

I'd like to see the solution myself

hero (hero):

Some problems are strange, cruel, and unusual. This is one of them

OpenStudy (anonymous):

haha tell me about it. So there's no way to graph it on the calculator and somehow come to a conclusion based on that? This problem is supposed to take two mins

hero (hero):

Maybe it really is a typo and your teacher hasn't told you yet

OpenStudy (anonymous):

maybe you can help me with this one? Let f be the function given by f(x) = x^2 - 2x +3. The tangent line to the graph of f at x = 2 is used to approximate values of f(x). what is the greatest value of x for which the error resulting from this tangent line approx. is less than 0.5?

OpenStudy (anonymous):

maybe but I doubt it bc it is photocopied right from the packet

hero (hero):

Oh I hate those....ask amistre

OpenStudy (anonymous):

ok. thank you. I will post the solution to that other one here once I get it

hero (hero):

Great